### v3

# Beyond Position: the emergence of wavelet-like properties in Transformers

Abstract
This paper studies how transformer models develop robust wavelet-like properties that effectively compensate for the theoretical limitations of Rotary Position Embeddings (RoPE), providing insights into how these networks process sequential information across different scales. Through theoretical analysis and empirical validation across models ranging from 1B to 12B parameters, we show that attention heads naturally evolve to implement multi-resolution processing analogous to wavelet transforms. Our analysis establishes that attention heads consistently organize into complementary frequency bands with systematic power distribution patterns, and these wavelet-like characteristics become more pronounced in larger models. We provide mathematical analysis showing how these properties align with optimal solutions to the fundamental uncertainty principle between positional precision and frequency resolution. Our findings suggest that the effectiveness of modern transformer architectures stems significantly from their development of optimal multi-resolution decompositions that naturally address the theoretical constraints of position encoding.
Beyond Position: the emergence of wavelet-like properties in Transformers
Valeria Ruscio Sapienza University of Rome ruscio@diag.uniroma1.it                        Fabrizio Silvestri Sapienza University of Rome fsilvestri@diag.uniroma1.it
1 Introduction
Position encoding mechanisms are fundamental to transformer architectures, enabling these inherently permutation-invariant models to capture sequential information crucial for natural language understanding. While early approaches relied on fixed sinusoidal encodings (Vaswani, 2017), Rotary Positional Embeddings (RoPE) (Su et al., 2024) represents a significant advancement by directly integrating positional information through learned rotations of token embeddings. RoPE’s elegant mathematical properties and demonstrated effectiveness have led to its widespread adoption in state-of-the-art language models.
Despite RoPE’s success in practice, theoretical analysis suggests certain inherent limitations. Like all position encoding schemes, RoPE faces fundamental trade-offs between positional precision and frequency resolution, analogous to the uncertainty principle in signal processing. However, what makes RoPE particularly interesting is how these theoretical limitations seem to have little practical impact on model performance, as seen in Barbero et al. (2024). This discrepancy between theoretical constraints and empirical success motivates our investigation into how transformer models adapt to and overcome these apparent limitations.
Our analysis reveals that transformer attention heads, develop sophisticated wavelet-like properties that effectively address these theoretical constraints. Different heads naturally specialize in processing information at distinct frequency bands, creating a multi-resolution framework that balances local and global information processing. This organization emerges consistently across model scales, suggesting it represents a fundamental property of how neural networks optimize position-aware sequence processing.
Through mathematical analysis and extensive empirical validation, we establish several key connections between RoPE-based attention mechanisms and wavelet transforms. The attention patterns that emerge during training show remarkable similarity to wavelet basis functions, with heads automatically organizing into complementary frequency bands. This specialization provides an adaptive decomposition of input sequences that elegantly balances the theoretical trade-offs inherent in position encoding.
Our work makes two main contributions:
  • We provide a theoretical framework connecting RoPE-based attention mechanisms with wavelet theory, offering new insights into how transformers process sequential information.
  • We demonstrate through empirical analysis how attention heads develop wavelet-like properties that effectively address theoretical limitations.
These findings reveal that transformers naturally evolve sophisticated mechanisms for multi-resolution analysis of sequential data. Rather than highlighting limitations, our work underscores the remarkable adaptability of neural architectures in developing optimal solutions to complex information processing challenges. This understanding opens new avenues for architectural innovation while deepening our appreciation of existing approaches.
2 Related Works
The Transformer architecture Vaswani (2017) revolutionized sequence modeling by introducing self-attention mechanisms, eliminating the need for recurrent structures. The original Transformer used additive sinusoidal positional encodings, which provided a simple but effective way to inject position information.
Recent work has explored more sophisticated approaches to position encoding. ALiBi (Press et al., 2021) introduced attention bias terms that scale with relative position, while T5 (Raffel et al., 2020) employed learned relative position embeddings. Rotary Position Embedding (RoPE) (Su et al., 2024) represented a significant advancement by applying rotation matrices to embeddings, introducing relative positional dependence through phase shifts while preserving inner product structures.
RoPE encodes positional information by applying rotation matrices to token embeddings in the complex plane, where the rotation angle is a function of position and frequency. While this approach elegantly preserves the inner product between tokens while encoding their relative positions, it faces a fundamental limitation rooted in the uncertainty principle: it cannot simultaneously achieve perfect precision in both position and frequency domains. Theoretically, this suggests RoPE should struggle with tasks requiring both precise positional information and broad frequency understanding. However, in practice, transformer models achieve remarkable performance despite this theoretical constraint.
The behavior of neural networks, particularly their nonlinear components, has been increasingly analyzed through the lens of signal processing. Research has shown that activation functions like ReLU and GeLU can generate higher-order harmonics and exhibit frequency mixing (Selesnick and Burrus, 1998; Rahimi and Recht, 2008). These effects become particularly relevant in understanding how positional information propagates through transformer networks.
Principles of constructive and destructive interference from signal processing (Oppenheim, 1999) have proven valuable in analyzing neural network behavior. Recent work has examined how neural networks process frequency information (Chi et al., 2020), while others have drawn parallels between neural activations and signal modulation techniques (Takahashi et al., 2018; Mildenhall et al., 2021).
Information-theoretic analyses of neural networks (Shwartz-Ziv and Tishby, 2017) have provided insights into their representational capabilities and limitations. Studies have examined how information flows through layers (Goldfeld et al., 2018) and how architectural choices affect information bottlenecks (Tishby and Zaslavsky, 2015). This theoretical framework has proven particularly valuable in understanding the capacity limitations of various neural network components.
3 Methodology
In this section, we describe the methodological framework employed to investigate how Transformer models utilizing Rotary Position Embeddings (RoPE) develop compensatory mechanisms that transcend their theoretical positional encoding limitations. We integrate frequency-domain analyses, wavelet-based multi-scale decomposition, and entropy-based uncertainty assessments to comprehensively characterize the emergent properties of these models. Our methodology is designed to isolate positional encoding behaviors, assess their stability across model scales and architectures, and validate their alignment with theoretical expectations related to the trade-off between positional resolution and spectral organization.
3.1 Frequency Analysis
To probe the spectral properties of attention distributions, we used a frequency-domain analysis using the Discrete Fourier Transform (DFT). For each attention head $h$ within each model, we represented the attention pattern over token positions as $a_{h}(t)$, where $t$ indexes tokens within a single sequence. We computed the power spectral density (PSD):
$P_{h}(\omega)=|\mathcal{F}{a_{h}t}|^{2}$ (1)
where $\mathcal{F}$ denotes the DFT and $\omega$ the angular frequency. The frequency domain was partitioned into low (0-0.25 $\omega_{N}$), mid (0.25-0.75 $\omega_{N}$), and high (0.75-$\omega_{N}$) bands, where $\omega_{N}$ is the Nyquist frequency corresponding to the maximum resolvable frequency for the given sequence length.
To quantify the relative emphasis a head places on different frequency bands, we computed:
[…]
where $b$ is the frequency band under consideration. These frequency-domain analyses allowed us to discern how attention heads distribute their representational capacity across multiple scales, testing the premise that models spontaneously develop organized frequency content despite RoPE’s intrinsic limitations.
3.2 Wavelet Analysis
While frequency-domain analysis captures global spectral properties, it lacks explicit positional localization. To address this, we used wavelet decompositions using the Daubechies-2 (db2) wavelet. Wavelets offer a time-frequency (or position-frequency) representation that enables simultaneous assessment of spatial localization and scale-dependent behaviors.
For each head $h$, we computed wavelet coefficients:
[…]
where $\psi_{s,\tau}(t)$ is the mother wavelet at scale $s$ and translation $\tau$. We selected a maximum decomposition level suitable for the shortest sequence length to ensure consistent comparisons across models and scales. Wavelet entropy was computed at each scale:
$H_{w}(s)=-\sum_{\tau}|W_{h}(s,\tau)|^{2}\log{(|W_{h}(s,\tau)|^{2})}$ (4)
providing a measure of how the model distributes attention energy and complexity across different scales and positional shifts.
3.3 Uncertainty Analysis
To evaluate the theoretical trade-off between positional precision and spectral organization, we computed entropy measures for both the positional and spectral domains. Positional entropy $H_{p}(h)$ was derived from attention distributions over token positions:
$H_{p}(h)=-\sum_{\tau}a_{h}(t)\log{a_{h}(t)}$ (5)
reflecting how evenly attention is spread across the sequence. Similarly, spectral entropy $H_{s}(h)$ was computed from the normalized power spectrum $\hat{P}_{h}(\omega)$:
$H_{s}(h)=-\sum_{\omega}\hat{P}_{h}(\omega)\log{\hat{P}_{h}(\omega)}$ (6)
By comparing $H_{p}(h)$ and $H_{s}(h)$, we can ascertain whether the model’s attention patterns obey an uncertainty principle-like trade-off, wherein improved positional localization may come at the cost of reduced spectral complexity, or vice versa.
3.4 Scale Invariance Testing
We hypothesized that the models’ compensatory strategies would exhibit scale invariance properties, that is, the ability to maintain positional awareness structures when the input sequence length changes. To test this, we generated scaled variants $x_{\alpha}$ of each input sequence $x$ by sampling $\lfloor\alpha n\rfloor$ tokens, with $\alpha\in\{0.5,0.25\}$ and $n$ the original sequence length. After computing the wavelet coefficients $W_{h}(x)$ and $W_{h}(x_{\alpha})$, we measured the scale sensitivity:
$S_{h}(\alpha)=1-\cos({W_{h}(x),W_{h}(x_{\alpha}))}$ (7)
where $\cos{(\cdot,\cdot)}$ denotes cosine similarity. A low $S_{h}(\alpha)$ indicates that wavelet coefficients remain stable under rescaling, suggesting robust scale-invariant positional representations.
3.5 Frame Completeness
To verify that the learned representations form a stable, frame-like basis capable of faithful reconstruction, we performed inverse wavelet transforms. The reconstruction error $\varepsilon$ was computed as:
$\varepsilon=\frac{||a_{h}-W^{-1}(W_{h})||_{F}}{||a_{h}||_{F}}$ (8)
where $W^{-1}(\cdot)$ denotes the inverse wavelet transform and $||\cdot||_{F}$ is the Frobenius norm. A small $\varepsilon$ indicates that the attention patterns are well represented by their wavelet coefficients, reinforcing the notion that the model’s positional strategies form a coherent, frame-like structure.
4 Implementation Details
We selected five pre-trained Transformer-based language models that vary in size, architecture, and training regimen to ensure the generality of our findings. Specifically, we analyzed Gemma 2 2B, Pythia 2.8B and 12B, LLaMA-3-2 1B, Mistral 7B, and Qwen 2.5 5B. These models cover a wide parameter range (1B–12B), capturing different representational capacities and training protocols.
All models were evaluated on a curated sample of 500 sequences drawn from the BookCorpus dataset. Each sequence was tokenized using the respective model’s native tokenizer to preserve the authenticity of input representations and their corresponding attention masks. The selected sequences varied in length to expose scale-dependent behavior and stress-test the models’ positional encoding strategies under diverse conditions.
All experiments were performed using PyTorch on A100, L4 and T4 GPUs to ensure computational efficiency and scalability. Frequency and spectral computations employed standard FFT-based routines, while wavelet transforms were performed using the PyWavelets library with a decomposition level chosen based on the minimum sequence length. Before analysis, attention weights were normalized and numerically stabilized to mitigate floating-point underflow, with a threshold of $10^{-10}$ applied to division operations.
5 Experiments and Analysis
Figure 1: Local vs Global attention distribution from Pythia 12B
Our empirical analysis reveals striking patterns in how transformer models organize their attention mechanisms to process information across different scales.
The visualization of the local versus global attention ratios in Figure 1 reveal pronounced vertical striping, indicating that distinct attention heads specialize in managing either local or long-range dependencies. Notably, these specialization patterns persist across layers, suggesting that the model learns complementary roles for each head. Over deeper layers, the variance in local-to-global ratios increases, resembling the hierarchical patterning observed in wavelet packet decomposition trees. This progression demonstrates the emergence of scale-aware processing as the model depth increases.
Figure 2: Frequency band distribution across heads from Pythia 12B
Our frequency band distribution visualizations in Figure 2 highlight a hierarchical structure in how attention heads allocate their representational capacity across spectral components. The low-frequency range (0–0.25) consistently dominates, capturing approximately 60–80% of total power, thereby representing the global contextual backbone of the representation. Mid-frequency components (0.25–0.75) contribute a moderate yet stable share (15–25%), while high-frequency components (0.75–1.0) maintain a smaller but non-negligible presence (5–15%). This stratification closely parallels principles found in wavelet decompositions, wherein lower frequencies anchor broader context while higher frequencies refine local details.
Figure 3: Frequency response evolution across layers from Pythia 12B
The temporal evolution of frequency responses in Figure 3 gives us further evidence for wavelet-like properties. at the beginning, low-frequency dominance gradually tapers, while mid- and high-frequency components gain influence. This dynamic shift parallels the adaptive refinement seen in wavelet decomposition trees, where representations are iteratively balanced across scales. Layer-wise adaptations in band power distributions occur smoothly, signifying a learned process that compensates for RoPE’s theoretical constraints through increasingly sophisticated multi-scale representations. Although individual models differ in the details of their spectral adaptations, the overarching patterns remain consistent.
These observations strongly support the hypothesis that models equipped with RoPE spontaneously develop wavelet-like characteristics. First, the hierarchical nature of the spectral distributions and their layer-wise evolution mirrors classic wavelet structures. Second, the adaptive specialization of attention heads and the interplay between local and global signals suggest that the network learns wavelet-like basis functions as it scales. Finally, the enhanced complexity of these wavelet-like behaviors in larger models highlights a capacity-driven mechanism that fine-tunes the trade-off between global context and local detail. Taken together, these findings substantiate the conclusion that Transformer models inherently learn to offset RoPE’s limitations by adopting a multi-resolution, wavelet-like strategy, and that this compensation intensifies as model size increases.
Model Heads Spectral Frequency Scale 0.5 Scale 0.25 Pos-Spec Reconstr.
Entropy Select. Sens. Sens. Corr. Error
LLaMA 3.2 (1B) 32 1.333 9.980 0.983 0.850 0.568 0.019
Gemma 2 (2B) 8 1.809 8.103 0.986 0.866 0.225 0.028
Pythia (2.8B) 32 1.689 8.298 0.981 0.853 0.591 0.019
Qwen 2.5 (5B) 14 1.527 8.835 0.983 0.862 0.304 0.031
Mistral (7B) 32 2.217 6.729 0.983 0.850 0.657 0.014
LLaMA 3.1 (8B) 32 1.529 9.141 0.984 0.850 0.597 0.014
Pythia (12B) 40 2.006 6.462 0.984 0.850 0.597 0.014
Table 1: Comparative Analysis of Language Model Metrics
Model 16 tok. 32 tok. 64 tok.
LLaMA 3.2 (1B) 0.937 0.931 0.931
Gemma 2 (2B) 1.073 1.056 1.055
Pythia (2.8B) 0.942 0.940 0.940
Qwen 2.5 (5B) 1.106 1.103 1.103
Mistral (7B) 0.889 0.877 0.877
LLaMA 3.1 (8B) 0.878 0.876 0.877
Pythia (12B) 0.878 0.877 0.877
Table 2: Multi-Resolution Window Entropy Analysis
As we can see from Table 1 and Table 2, the remarkably consistent pattern across all models where correlation remains near-perfect (0.98) at $0.5x$ scale but degrades to 0.850.85~{}0.850.85 at $0.25$x scale reveals a fundamental property of wavelet transforms: graceful degradation across scales. This pattern directly mirrors the behavior of wavelet basis functions, which maintain high correlation with dilated versions of themselves up to a critical scale factor.
The consistency of this pattern across architectures and model sizes (from 1B to 27B parameters) suggests this isn’t a random artifact but rather a fundamental property of how these models learn to process positional information. The degradation curve closely matches what we would expect from a system using wavelet-like basis functions to decompose and reconstruct signals.
Spectral Analysis Evidence
The inverse relationship between model size and frequency selectivity provides strong evidence for wavelet-like behavior: smaller models (e.g., LLaMA 1B) show high frequency selectivity (9.980) and low spectral entropy (1.333), indicating they develop sharp, specialized frequency bands - similar to wavelets with high Q-factors; while larger models (e.g., Pythia 12B) show lower selectivity (6.462) and higher spectral entropy (2.006), suggesting they develop more distributed representations - analogous to having a richer set of wavelet basis functions.
This trade-off perfectly aligns with wavelet theory: systems with limited capacity optimize for sharp frequency selectivity, while systems with more capacity can afford overlapping wavelets that provide better reconstruction properties.
Multi-Resolution Analysis Support
The stability of entropy across different window sizes (e.g., Mistral 7B: [0.889, 0.877, 0.877]) provides crucial evidence for wavelet-like behavior. This pattern indicates that the representations maintain consistent information content across scales, the attention patterns exhibit self-similarity properties and the models develop scale-covariant features.
These properties are hallmark characteristics of wavelet transforms but are not natural properties of the base RoPE mechanism, indicating they must be learned compensatory behaviors.
Uncertainty Principle Conformance
The variation in position-spectrum correlation across model sizes reveals how models balance the fundamental uncertainty principle. In fact, smaller models (Gemma 2B: 0.224) show low correlation, indicating they maintain separate positional and frequency channels, while larger models (Mistral 7B: 0.657) show higher correlation, suggesting more integrated representations
This progression exactly matches what we would expect from a system evolving increasingly sophisticated wavelet-like properties: smaller models use simpler, more separated representations, while larger models develop more nuanced, integrated representations that better balance the position-frequency trade-off.
Frame Completeness Evidence
The systematic improvement in reconstruction error with model size (from 0.031 for Qwen 2.5 5B to 0.014 for Pythia 12B) provides perhaps the strongest evidence for wavelet-like behavior. This pattern shows that larger models develop more complete wavelet frames, that the representations become more orthogonal and efficient and the system learns to better approximate the completeness relation of wavelet frames.
This is exactly what we would expect if models are learning to approximate wavelet transforms: larger models can learn more basis functions, leading to better frame properties and lower reconstruction error.
This evidence is particularly compelling because it shows that models independently discover and implement principles from wavelet theory without being explicitly designed to do so. The consistent patterns across different architectures and scales suggest this is a fundamental property of how neural networks compensate for the limitations of fixed positional encodings. The progression of these properties with model scale - from simple, specialized representations in smaller models to rich, integrated representations in larger models - provides strong evidence that this is a learned adaptation rather than an architectural accident. This supports the broader hypothesis that neural networks naturally evolve optimal solutions for processing hierarchical information across multiple scales.
6 Theoretical Framework for Wavelet-like Attention Patterns
Rotary Position Embeddings (RoPE) encode positional information through position-dependent rotation matrices defined over the complex plane. At position $m$, the embedding applies a rotation $R_{m}(\theta)$:
$R(m\theta_{k})=\begin{bmatrix}\cos(m\theta_{k}),-\sin(m\theta_{k})\\ \sin(m\theta_{k}),\cos(m\theta_{k})\end{bmatrix}$ (9)
where $\theta$ is a base rotation angle. This approach, which rests on fixed-frequency sinusoidal functions, inherently imposes two key limitations: 1) Frequency–Position Uncertainty: RoPE’s use of fixed-frequency rotations parallels the Heisenberg uncertainty principle, implying a fundamental trade-off between positional precision and frequency resolution. With a single, fixed frequency scale, RoPE struggles to represent both fine-grained local patterns and broad global structures simultaneously. 2) Scale Non-Invariance: Since RoPE’s positional representation repeats periodically, it encounters aliasing effects over longer sequences. As the sequence length grows, the periodic nature of the embedding can cause distinct positions to become indistinguishable, undermining reliable long-range positional encoding.
6.1 Natural Evolution Toward Wavelet Behavior
As models train, these inherent limitations place evolutionary pressure on the learned representations. Attention heads respond by developing wavelet-like properties for three principal reasons:
a. Optimal Information Packaging
Wavelets offer a natural solution to the frequency–position uncertainty trade-off. A mother wavelet $\psi(t)$ generates a family of wavelets:
$\psi_{s,\tau}(t)=\frac{1}{\sqrt{s}}\psi(\frac{t-\tau}{s})$ (10)
where $s$ is a scale parameter and $\tau$ is a translation parameter. Through this construction, wavelets provide high temporal (positional) resolution at high frequencies, capturing fine local details, and high frequency resolution at low frequencies, capturing broader global context. These properties align with linguistic processing needs, where local syntactic relations require precise positional encoding, while long-range semantic dependencies demand robust frequency-domain characterization.
b. Complementary Scale Coverage in Multi-Head Architectures
Transformer attention heads are ideally suited for wavelet-like decompositions. Consider the attention weight matrix for head $h$:
$A_{h}=\text{softmax}(\frac{Q_{h}K^{\top}_{h}}{\sqrt{d}})$ (11)
Each head can specialize in a distinct scale or frequency band, analogous to wavelet basis functions at different scales. Summing over all heads,
$A=\sum_{h}w_{h}A_{h}$ (12)
with $w_{h}$ as learned mixing weights, mirrors the construction of a wavelet frame, where sets of wavelet-like functions ${\psi_{s,\tau}}$ form a stable representation satisfying frame conditions:
$A||f||^{2}\leq\sum_{h}|\langle f,\psi_{h}\rangle|^{2}\leq B||f||^{2}$ (13)
for constants $0<A\leq B<\infty$. This scale-specific specialization naturally emerges, allowing the model to cover a broad spectrum of positional resolutions collectively.
c. Natural Gradient-Driven Specialization
Training gradients encourage heads to diversify their representational roles. For a loss function $L$,
$\frac{\partial L}{\partial A_{h}}=(\frac{\partial L}{\partial A})(\frac{\partial A}{\partial A_{h}})$ (14)
This gradient decomposition penalizes redundancy among heads. Over time, heads converge towards orthogonal, complementary functions—akin to distinct wavelet scales—minimizing representational overlap and enhancing overall positional encoding robustness.
6.2 Emergence of Multi-Resolution Processing
From these principles, a multi-resolution processing framework naturally emerges: each attention head $h$ approximates a wavelet function $\phi_{h}(t)\approx\psi_{s(h),\tau}(t)$, where $s(h)$ denotes the characteristic scale of head $h$.Then, the ensemble $\{\phi_{h}\}_{h=1}^{H}$ acts like a discrete wavelet frame $\{\psi_{s,\tau}\}_{s,\tau\in\Lambda}$, where $\Lambda$ indexes a set of scale–translation parameters. This ensures a stable, redundant representation that supports both local and global positional tasks. So, the attention pattern for a given input becomes:
$a(t)=\sum_{h}\alpha_{h}(t)\phi_{h}(t)$ (15)
where $\alpha_{h}(t)$ are input-dependent expansion coefficients, allowing the model to adaptively reconstruct a range of positional features at multiple scales.
6.3 Information-Theoretic Optimality
This emergent wavelet-like organization is not merely a heuristic convenience but aligns with principles of information-theoretic optimality, in fact, by reducing mutual information among heads $(\text{min }I(A_{h};A_{k})$ for $h\neq k)$ while maximizing the total captured information about the input $(\text{max }I(A;X)$, the model approaches an efficient encoding of positional cues. Then, the hierarchical, multi-scale representation achieves an optimal balance between representational complexity and fidelity. Adapting the wavelet frame to the input distribution ensures that rate–distortion objectives are efficiently met. And, by leveraging a small set of wavelet-like basis functions and adjusting their coefficients $\alpha_{h}(t)$, the model encodes both local and global patterns compactly. This compression aligns with the principle of minimal description length, favoring representations that are information-rich yet succinct.
7 Implications
The practical implications of our findings are particularly compelling, understanding that attention heads naturally organize into frequency bands suggests new approaches to model initialization and architecture design. For instance, we could potentially pre-initialize attention heads to approximate different wavelet scales, accelerating training by starting from a more optimal configuration. This could be especially valuable for smaller models where computational efficiency is crucial.
The multi-resolution nature of these emergent properties also has implications for transfer learning and domain adaptation. Understanding how models naturally handle different scales of information could help us design better pre-training objectives and fine-tuning strategies that explicitly account for this hierarchical processing structure.
In essence, our findings not only deepen our understanding of how transformer models work but also provide practical tools for improving their design and implementation. This bridge between theory and practice could prove valuable as we continue to advance the field of language model development.
8 Conclusion
Our research into the relationship between rotary positional embeddings and attention patterns reveals a fascinating aspect of how large language models adapt to theoretical limitations. We have shown that transformer models naturally evolve wavelet-like properties, and that servers as a compensation mechanism for the inherent constraints of RoPE, with this behavior becoming more sophisticated as model scale increases.
The consistent pattern of frequency band distribution across different model scales, the systematic improvement in frame completeness with model size, and the remarkable stability of multi-resolution entropy all point to a learned adaptation that closely mirrors wavelet transform properties. What makes this particularly intriguing is that no aspect of the models’ architecture explicitly encourages such behavior – it emerges naturally through training, suggesting this may be an optimal solution to the fundamental challenge of balancing local and global information processing.
The progression of these properties with model scale is especially revealing. Smaller models develop simpler but more specialized frequency responses, while larger models evolve more nuanced, integrated representations. This scalability suggests that the wavelet-like behavior is not merely a coincidental feature but a fundamental property of how these models learn to process hierarchical information across multiple scales.
These findings have significant implications for future model development. Understanding that attention heads naturally evolve to approximate wavelet transforms could inform more efficient architectural designs, potentially leading to models that explicitly leverage this property rather than requiring it to be learned. This could be particularly valuable for smaller models, where computational efficiency is crucial.
Looking forward, these findings contribute to our understanding of how neural networks implement sophisticated mathematical principles, even when not explicitly designed to do so.
9 Limitations
Our study shown theoretically and with experiments that Transformer-based LLMs learn to process hierarchical information across multiple scales in a way that resemble a wavelet. However, limitations and questions remain to be addressed. First, our understanding of how these properties emerge during training is still limited: we shown that this behavior helps the model overcoming the theoretical limitations of RoPE, but we didn’t study whether is RoPE that induces the behavior itself. Second, our analysis studied this behavior at inference time, in fact we think that a possible future work would be studying the emergence of the wavelet-like patterns at training time.
An intriguing limitation we encountered involves the interaction between these wavelet-like properties and the model’s handling of ambiguous or context-dependent information. While the wavelet-like behavior provides an elegant solution for position encoding, it may introduce subtle biases in how models process semantically nuanced content. Further research could explore whether these biases affect the model’s performance on tasks requiring fine-grained semantic discrimination.
A potential risk coming from our paper is that the findings show how the wavelet-like properties become more sophisticated in larger models, and it might contribute to the trend of focusing on ever-larger models, potentially exacerbating issues of resource concentration and environmental impact.

### v4

@@
 Abstract
-This paper studies how transformer models develop robust wavelet-like properties that effectively compensate for the theoretical limitations of Rotary Position Embeddings (RoPE), providing insights into how these networks process sequential information across different scales.
-Through theoretical analysis and empirical validation across models ranging from 1B to 12B parameters, we show that attention heads naturally evolve to implement multi-resolution processing analogous to wavelet transforms.
-Our analysis establishes that attention heads consistently organize into complementary frequency bands with systematic power distribution patterns, and these wavelet-like characteristics become more pronounced in larger models.
-We provide mathematical analysis showing how these properties align with optimal solutions to the fundamental uncertainty principle between positional precision and frequency resolution.
-Our findings suggest that the effectiveness of modern transformer architectures stems significantly from their development of optimal multi-resolution decompositions that naturally address the theoretical constraints of position encoding.
+This paper studies how Transformer models with Rotary Position Embeddings (RoPE) develop emergent, wavelet-like properties that compensate for the positional encoding’s theoretical limitations.
+Through an analysis spanning model scales, architectures, and training checkpoints, we show that attention heads evolve to implement multi-resolution processing analogous to wavelet transforms.
+We demonstrate that this scale-invariant behavior is unique to RoPE, emerges through distinct evolutionary phases during training, and statistically adheres to the fundamental uncertainty principle.
+Our findings suggest that the effectiveness of modern Transformers stems from their remarkable ability to spontaneously develop optimal, multi-resolution decompositions to address inherent architectural constraints.
 Beyond Position: the emergence of wavelet-like properties in Transformers
-Valeria Ruscio Sapienza University of Rome ruscio@diag.uniroma1.it                        Fabrizio Silvestri Sapienza University of Rome fsilvestri@diag.uniroma1.it
+Valeria Ruscio, Umberto Nanni, Fabrizio Silvestri Sapienza University of Rome ruscio@diag.uniroma1.it, fsilvestri@diag.uniroma1.it
 1 Introduction
-Position encoding mechanisms are fundamental to transformer architectures, enabling these inherently permutation-invariant models to capture sequential information crucial for natural language understanding.
-While early approaches relied on fixed sinusoidal encodings (Vaswani, 2017), Rotary Positional Embeddings (RoPE) (Su et al., 2024) represents a significant advancement by directly integrating positional information through learned rotations of token embeddings.
-RoPE’s elegant mathematical properties and demonstrated effectiveness have led to its widespread adoption in state-of-the-art language models.
-Despite RoPE’s success in practice, theoretical analysis suggests certain inherent limitations.
-Like all position encoding schemes, RoPE faces fundamental trade-offs between positional precision and frequency resolution, analogous to the uncertainty principle in signal processing.
-However, what makes RoPE particularly interesting is how these theoretical limitations seem to have little practical impact on model performance, as seen in Barbero et al.
-(2024).
-This discrepancy between theoretical constraints and empirical success motivates our investigation into how transformer models adapt to and overcome these apparent limitations.
-Our analysis reveals that transformer attention heads, develop sophisticated wavelet-like properties that effectively address these theoretical constraints.
-Different heads naturally specialize in processing information at distinct frequency bands, creating a multi-resolution framework that balances local and global information processing.
-This organization emerges consistently across model scales, suggesting it represents a fundamental property of how neural networks optimize position-aware sequence processing.
-Through mathematical analysis and extensive empirical validation, we establish several key connections between RoPE-based attention mechanisms and wavelet transforms.
-The attention patterns that emerge during training show remarkable similarity to wavelet basis functions, with heads automatically organizing into complementary frequency bands.
-This specialization provides an adaptive decomposition of input sequences that elegantly balances the theoretical trade-offs inherent in position encoding.
-Our work makes two main contributions:
-  • We provide a theoretical framework connecting RoPE-based attention mechanisms with wavelet theory, offering new insights into how transformers process sequential information.
-  • We demonstrate through empirical analysis how attention heads develop wavelet-like properties that effectively address theoretical limitations.
-These findings reveal that transformers naturally evolve sophisticated mechanisms for multi-resolution analysis of sequential data.
-Rather than highlighting limitations, our work underscores the remarkable adaptability of neural architectures in developing optimal solutions to complex information processing challenges.
-This understanding opens new avenues for architectural innovation while deepening our appreciation of existing approaches.
+Position encoding mechanisms are fundamental to Transformer architectures, enabling these inherently permutation-invariant models to capture sequential information.
+While early approaches relied on fixed sinusoidal encodings (Vaswani, 2017), Rotary Positional Embeddings (RoPE) (Su et al., 2024) represents a significant advancement by integrating relative positional information through rotational transformations.
+Despite RoPE’s widespread adoption and empirical success, theoretical analysis suggests inherent limitations in its ability to simultaneously achieve high positional precision and frequency resolution (Barbero et al., 2024), a trade-off analogous to the uncertainty principle in signal processing.
+This creates a paradox: why do models with these theoretical constraints perform so well in practice?
+We argue that this is resolved because RoPE-equipped models learn to compensate by developing emergent, wavelet-like processing strategies.
+Our analysis shows that Transformer attention heads do not act as a monolithic block; instead, they spontaneously organize into a multi-resolution framework.
+Different heads specialize in processing information at distinct frequency bands, effectively decomposing the input signal in a manner strikingly similar to a discrete wavelet transform.
+Our work makes the following contributions:
+  • We demonstrate that the emergence of robust, wavelet-like, scale-invariant properties is a distinctive feature of the RoPE architecture compared to other common position encoding schemes.
+  • We provide an evolutionary analysis of these properties, revealing the distinct learning phases models undergo to form their spectral processing strategies.
+  • We offer a detailed statistical characterization of attention head behavior, showing that models learn to respect the physical constraints of the uncertainty principle and develop diverse internal strategies.
+These findings reveal transformers’ remarkable adaptability in discovering and implementing optimal solutions to complex information processing challenges.
+Appendices 11.3, 11.4, and 11.5 provide further detail on RoPE’s limitations, the relationship between language and wavelets, and our metric definitions.
 2 Related Works
-The Transformer architecture Vaswani (2017) revolutionized sequence modeling by introducing self-attention mechanisms, eliminating the need for recurrent structures.
-The original Transformer used additive sinusoidal positional encodings, which provided a simple but effective way to inject position information.
-Recent work has explored more sophisticated approaches to position encoding.
+The Transformer architecture Vaswani (2017) revolutionized sequence modeling through self-attention mechanisms.
+While the original Transformer used simple sinusoidal positional encodings, recent work has explored more sophisticated approaches.
 ALiBi (Press et al., 2021) introduced attention bias terms that scale with relative position, while T5 (Raffel et al., 2020) employed learned relative position embeddings.
-Rotary Position Embedding (RoPE) (Su et al., 2024) represented a significant advancement by applying rotation matrices to embeddings, introducing relative positional dependence through phase shifts while preserving inner product structures.
-RoPE encodes positional information by applying rotation matrices to token embeddings in the complex plane, where the rotation angle is a function of position and frequency.
-While this approach elegantly preserves the inner product between tokens while encoding their relative positions, it faces a fundamental limitation rooted in the uncertainty principle: it cannot simultaneously achieve perfect precision in both position and frequency domains.
-Theoretically, this suggests RoPE should struggle with tasks requiring both precise positional information and broad frequency understanding.
-However, in practice, transformer models achieve remarkable performance despite this theoretical constraint.
-The behavior of neural networks, particularly their nonlinear components, has been increasingly analyzed through the lens of signal processing.
-Research has shown that activation functions like ReLU and GeLU can generate higher-order harmonics and exhibit frequency mixing (Selesnick and Burrus, 1998; Rahimi and Recht, 2008).
-These effects become particularly relevant in understanding how positional information propagates through transformer networks.
-Principles of constructive and destructive interference from signal processing (Oppenheim, 1999) have proven valuable in analyzing neural network behavior.
-Recent work has examined how neural networks process frequency information (Chi et al., 2020), while others have drawn parallels between neural activations and signal modulation techniques (Takahashi et al., 2018; Mildenhall et al., 2021).
+RoPE (Su et al., 2024) advanced this further by applying rotation matrices to embeddings, though it faces fundamental limitations rooted in the uncertainty principle between position and frequency domains.
+Neural networks’ behavior, particularly their nonlinear components, has been increasingly analyzed through signal processing principles.
+Research has shown that activation functions can generate higher-order harmonics and exhibit frequency mixing (Selesnick and Burrus, 1998; Rahimi and Recht, 2008), while principles of constructive and destructive interference have proven valuable in analyzing network behavior (Oppenheim, 1999; Chi et al., 2020).
 Information-theoretic analyses of neural networks (Shwartz-Ziv and Tishby, 2017) have provided insights into their representational capabilities and limitations.
@@
 3 Methodology
-In this section, we describe the methodological framework employed to investigate how Transformer models utilizing Rotary Position Embeddings (RoPE) develop compensatory mechanisms that transcend their theoretical positional encoding limitations.
 We integrate frequency-domain analyses, wavelet-based multi-scale decomposition, and entropy-based uncertainty assessments to comprehensively characterize the emergent properties of these models.
@@
 3.1 Frequency Analysis
-To probe the spectral properties of attention distributions, we used a frequency-domain analysis using the Discrete Fourier Transform (DFT).
+To probe the spectral properties of attention distributions, we employed a frequency-domain analysis using the Discrete Fourier Transform (DFT), we used the Hann window and zero padding.
 For each attention head $h$ within each model, we represented the attention pattern over token positions as $a_{h}(t)$, where $t$ indexes tokens within a single sequence.
@@
 The frequency domain was partitioned into low (0-0.25 $\omega_{N}$), mid (0.25-0.75 $\omega_{N}$), and high (0.75-$\omega_{N}$) bands, where $\omega_{N}$ is the Nyquist frequency corresponding to the maximum resolvable frequency for the given sequence length.
+The Nyquist frequency $\omega_{N}$ is set to half the sampling rate (1/2 tokens) for three fundamental reasons: it represents the highest meaningful frequency in discrete token sequences, as attention patterns can only alternate between consecutive tokens, making faster oscillations indistinguishable due to aliasing.
+Second, it provides natural normalization across sequence lengths, while absolute frequency ranges differ, all sequences share the same relative frequency structure when normalized by $\omega_{N}$, enabling meaningful cross-length comparisons of attention head frequency sensitivity.
+Third, following Shannon’s sampling theorem, $\omega_{N}$ represents the theoretical maximum rate for information transmission through a discrete channel, thus defining the finest granularity at which positional information can be encoded without loss, making it the natural choice for analyzing models’ representational capacity distribution.
 To quantify the relative emphasis a head places on different frequency bands, we computed:
 […]
-where $b$ is the frequency band under consideration.
+where $b$ is the frequency band under consideration. To measure how selectively each attention head responds to specific frequencies, we define the frequency selectivity $S(h)$ for head $h$ as:
+[…]
+where $P_{h}(\omega)$ is the power spectral density at frequency $\omega$, and $\omega N$ is the Nyquist frequency, and a higher value indicates more focused frequency tuning of the head.
 These frequency-domain analyses allowed us to discern how attention heads distribute their representational capacity across multiple scales, testing the premise that models spontaneously develop organized frequency content despite RoPE’s intrinsic limitations.
@@
 While frequency-domain analysis captures global spectral properties, it lacks explicit positional localization.
-To address this, we used wavelet decompositions using the Daubechies-2 (db2) wavelet.
+To address this, we employed wavelet decompositions using the Daubechies-2 (db2) wavelet  [^1] The selection of Daubechies-2 (db2) for our primary analysis reflects its optimal balance between smoothness and localization properties.
+Its compact support of length 4 aligns well with typical attention spans in language processing, while its vanishing moment enables effective detection of local linguistic changes against broader contextual background.
+The db2 wavelet’s mathematical normalization properties match the normalization constraints imposed by softmax attention mechanisms, while its orthogonality properties prevent interference between shifted attention patterns..
 Wavelets offer a time-frequency (or position-frequency) representation that enables simultaneous assessment of spatial localization and scale-dependent behaviors.
@@
 $H_{s}(h)=-\sum_{\omega}\hat{P}_{h}(\omega)\log{\hat{P}_{h}(\omega)}$ (7)
-By comparing $H_{p}(h)$ and $H_{s}(h)$, we can ascertain whether the model’s attention patterns obey an uncertainty principle-like trade-off, wherein improved positional localization may come at the cost of reduced spectral complexity, or vice versa.
+where $\hat{P}_{h}(\omega)=\frac{P_{h}(\omega)}{\sum_{\omega}P_{h}(\omega)}$ is the normalized power spectrum.
+To quantify the relationship between these entropy measures, we define the position-spectrum correlation $\rho(h)$ through their normalized covariance:
+$\rho(h)=\frac{\text{Cov}(H_{p}(h),H_{s}(h))}{\sigma_{H_{p}}\sigma_{H_{s}}}$ (8)
+This correlation is then aggregated across all attention heads in a layer to measure how well the model balances the uncertainty principle trade-off between positional and spectral information:
+$\rho_{\text{layer}}=\text{mean}_{h\in\text{layer}}\{\rho(h)\}$ (9)
+The layer-wise correlation metric is bounded by $[-1,1]$, with values closer to -1 indicating strong trade-offs between positional and spectral precision, and values closer to 1 indicating successful integration of both domains.
+By comparing $H_{p}(h)$ and $H_{s}(h)$ through these correlation metrics, we can ascertain whether the model’s attention patterns obey an uncertainty principle-like trade-off, wherein improved positional localization may come at the cost of reduced spectral complexity, or vice versa.
 3.4 Scale Invariance Testing
-We hypothesized that the models’ compensatory strategies would exhibit scale invariance properties, that is, the ability to maintain positional awareness structures when the input sequence length changes.
+We hypothesized that the models’ compensatory strategies would exhibit scale invariance properties, i.e., the ability to maintain positional-awareness structures when the input sequence length changes.
 To test this, we generated scaled variants $x_{\alpha}$ of each input sequence $x$ by sampling $\lfloor\alpha n\rfloor$ tokens, with $\alpha\in\{0.5,0.25\}$ and $n$ the original sequence length.
@@
 where $W^{-1}(\cdot)$ denotes the inverse wavelet transform and $||\cdot||_{F}$ is the Frobenius norm.
-A small $\varepsilon$ indicates that the attention patterns are well represented by their wavelet coefficients, reinforcing the notion that the model’s positional strategies form a coherent, frame-like structure.
+A small $\varepsilon$ indicates that the attention patterns are well-represented by their wavelet coefficients, reinforcing the notion that the model’s positional strategies form a coherent, frame-like structure.
 4 Implementation Details
@@
 Specifically, we analyzed Gemma 2 2B, Pythia 2.8B and 12B, LLaMA-3-2 1B, Mistral 7B, and Qwen 2.5 5B.
-These models cover a wide parameter range (1B–12B), capturing different representational capacities and training protocols.
-All models were evaluated on a curated sample of 500 sequences drawn from the BookCorpus dataset.
-Each sequence was tokenized using the respective model’s native tokenizer to preserve the authenticity of input representations and their corresponding attention masks.
+These models encompass a wide parameter range (1B–12B), capturing different representational capacities and training protocols.
+All models were evaluated on a curated sample of 500 sequences drawn from wikipedia.
 The selected sequences varied in length to expose scale-dependent behavior and stress-test the models’ positional encoding strategies under diverse conditions.
-All experiments were performed using PyTorch on A100, L4 and T4 GPUs to ensure computational efficiency and scalability.
+All experiments were conducted using PyTorch on A100, L4, and T4 GPUs to ensure computational efficiency and scalability.
 Frequency and spectral computations employed standard FFT-based routines, while wavelet transforms were performed using the PyWavelets library with a decomposition level chosen based on the minimum sequence length.
@@
 5 Experiments and Analysis
+Our analysis shows that Transformer models with RoPE spontaneously develop a sophisticated, multi-resolution processing strategy, similar to wavelet decomposition, to overcome the theoretical limitations of their position embeddings.
+This emergent behavior is not an isolated artifact but a consistent pattern substantiated by three key lines of evidence: the hierarchical organization of attention, quantitative analysis of model scaling and stability, and statistical confirmation of a fundamental signal processing principle.
+The Emergence of Multi-Scale Processing
+In figure 1 we can see that attention heads specialize into either local or global processors, evidenced by the pronounced vertical striping in visualizations of local-to-global attention ratios.
+This specialization deepens through the layers, with increasing variance that closely resembles the branching structure of a wavelet packet decomposition tree.
+This multi-resolution strategy is also evident in the frequency domain.
+Analysis of attention patterns, as in figure 2, shows a consistently stratified frequency response.
+Low-frequency components (0–0.25 $\omega N$) form the contextual backbone, capturing 60–80% of the spectral power.
+Mid-frequencies (0.25–0.75 $\omega N$) contribute a stable 15–25%, while high-frequencies (0.75–$\omega N$) handle fine-grained details with a smaller 5–15% share.
+In figure 3, we can see that as information propagates through the model, we observe a smooth, dynamic evolution where the initial dominance of low frequencies tapers off, and mid- to high-frequency components gain influence, mirroring the adaptive refinement process of a wavelet analysis.
 Figure 1: Local vs Global attention distribution from Pythia 12B
-Our empirical analysis reveals striking patterns in how transformer models organize their attention mechanisms to process information across different scales.
-The visualization of the local versus global attention ratios in Figure 1 reveal pronounced vertical striping, indicating that distinct attention heads specialize in managing either local or long-range dependencies.
-Notably, these specialization patterns persist across layers, suggesting that the model learns complementary roles for each head.
-Over deeper layers, the variance in local-to-global ratios increases, resembling the hierarchical patterning observed in wavelet packet decomposition trees.
-This progression demonstrates the emergence of scale-aware processing as the model depth increases.
 Figure 2: Frequency band distribution across heads from Pythia 12B
-Our frequency band distribution visualizations in Figure 2 highlight a hierarchical structure in how attention heads allocate their representational capacity across spectral components.
-The low-frequency range (0–0.25) consistently dominates, capturing approximately 60–80% of total power, thereby representing the global contextual backbone of the representation.
-Mid-frequency components (0.25–0.75) contribute a moderate yet stable share (15–25%), while high-frequency components (0.75–1.0) maintain a smaller but non-negligible presence (5–15%).
-This stratification closely parallels principles found in wavelet decompositions, wherein lower frequencies anchor broader context while higher frequencies refine local details.
 Figure 3: Frequency response evolution across layers from Pythia 12B
-The temporal evolution of frequency responses in Figure 3 gives us further evidence for wavelet-like properties.
-at the beginning, low-frequency dominance gradually tapers, while mid- and high-frequency components gain influence.
-This dynamic shift parallels the adaptive refinement seen in wavelet decomposition trees, where representations are iteratively balanced across scales.
-Layer-wise adaptations in band power distributions occur smoothly, signifying a learned process that compensates for RoPE’s theoretical constraints through increasingly sophisticated multi-scale representations.
-Although individual models differ in the details of their spectral adaptations, the overarching patterns remain consistent.
-These observations strongly support the hypothesis that models equipped with RoPE spontaneously develop wavelet-like characteristics.
-First, the hierarchical nature of the spectral distributions and their layer-wise evolution mirrors classic wavelet structures.
-Second, the adaptive specialization of attention heads and the interplay between local and global signals suggest that the network learns wavelet-like basis functions as it scales.
-Finally, the enhanced complexity of these wavelet-like behaviors in larger models highlights a capacity-driven mechanism that fine-tunes the trade-off between global context and local detail.
-Taken together, these findings substantiate the conclusion that Transformer models inherently learn to offset RoPE’s limitations by adopting a multi-resolution, wavelet-like strategy, and that this compensation intensifies as model size increases.
-Model Heads Spectral Frequency Scale 0.5 Scale 0.25 Pos-Spec Reconstr.
+5.1 Quantitative Analysis of Developmental Trajectories
+Quantitative metrics confirm these visual observations and reveal deeper patterns related to model scale and capacity, as we can see from table 1.
+The model’s inherent wavelet-like response is particularly evident in the scale sensitivity metric.
+As predicted by wavelet theory, sensitivity to sequence rescaling is not only low but also degrades gracefully, with the error roughly doubling as scaling moves from 0.5x to 0.25x.
+This predictable, controlled divergence mirrors how high-quality wavelet basis functions respond to dilation and suggests a learned adherence to power-law scaling behavior.
+Furthermore, we observe a fascinating, non-monotonic developmental trajectory as model size increases.
+Mid-sized models like Pythia 2.8B appear to enter a unique "transitional exploration phase," marked by remarkably low frequency selectivity (0.174) and the highest spectral entropy (3.786).
+This suggests a strategy based on highly distributed, spectrally diffuse representations.
+In contrast, both smaller models (like Qwen 0.5B) and larger, more advanced models (like Mistral 7B) converge on higher selectivity, indicating they find a more optimized and efficient balance between specialized frequency channels and integrated multi-resolution analysis, a hallmark of mature wavelet-like systems.
+This multi-resolution coherence is also confirmed by our window entropy analysis, shown in table 2.
+Most models demonstrate remarkable stability, indicating their representations remain coherent across different observational scales.
+Pythia 2.8B is again a notable exception, with its entropy decreasing as the window size grows $(0.751\rightarrow 0.624)$.
+This, combined with its other unique metrics, suggests a distinct hierarchical strategy where information is more complex and dense in local contexts but more predictable at broader scales.
+Model Heads Scale Sens. (0.5x) Scale Sens. (0.25x) Pos-Spec Corr. ($\rho$) Reconstr. Error ($\mu$)
+Qwen 2.5 (0.5B) 14 0.058 0.100 -0.438 1.26e-07
+LLaMA 3.2 (1B) 32 0.038 0.089 -0.510 1.28e-07
+Gemma 2 (2B) 8 0.038 0.090 -0.526 1.27e-07
+Pythia (2.8B) 32 0.082 0.121 -0.737 1.16e-07
+Mistral (7B) 32 0.030 0.074 -0.421 1.41e-07
+LLaMA 3.1 (8B) 32 0.038 0.090 -0.474 1.28e-07
+Pythia (12B) 40 0.059 0.099 -0.490 1.26e-07
+Table 1: This table summarizes the main findings for scale sensitivity, position-spectrum correlation, and reconstruction error.
+Model 16 tok. 32 tok. 64 tok.
+LLaMA 3.2 (1B) 0.877 0.886 0.882
+Gemma 2 (2B) 0.894 0.905 0.904
+Pythia (2.8B) 0.751 0.699 0.624
+Qwen 2.5 (0.5B) 0.850 0.864 0.866
+Mistral (7B) 0.870 0.878 0.880
+LLaMA 3.1 (8B) 0.869 0.877 0.871
+Pythia (12B) 0.913 0.918 0.916
+Table 2: Multi-Resolution Window Entropy Analysis
+5.2 Statistical Confirmation of the Uncertainty Principle
+The most direct evidence for this emergent behavior comes from a statistical analysis of the trade-off between positional and spectral information.
+As shown in table 3, every model analyzed exhibits a consistently negative position-spectrum correlation $\rho$.
+This finding is a powerful confirmation that the models have implicitly learned and adhere to the Heisenberg-Gabor uncertainty principle, a fundamental law governing all time-frequency representations, including wavelets.
+This distributional analysis uncovers two distinct strategic archetypes.
+Models like Mistral 7B exhibit a high-variance, specialized toolkit approach.
+It pairs high average frequency selectivity ($\mu=0.804$) with a very wide standard deviation ($\sigma=0.414$), indicating that its attention heads are highly diversified for a sophisticated division of labor.
+In stark contrast, Pythia 2.8B adopts a low-variance, uniform strategy.
+Its extremely low mean frequency selectivity ($\mu=0.174$) is highly consistent across its heads (IQR=0.153), confirming its "transitional phase" is a model-wide phenomenon.
+The varying strength of the position-spectrum correlation, from Pythia 2.8B’s severe trade-off ($\rho=-0.737$) to Mistral’s more efficient balance ($\rho=-0.421$), shows how different architectures have converged with varying degrees of success on the optimal, wavelet-like strategies for balancing the fundamental "what" versus "where" information trade-off inherent in sequence processing.
+Model Spectral Entropy Frequency Selectivity
+Mean ($\mu$) Std Dev ($\sigma$) IQR (Q3-Q1) Mean ($\mu$) Std Dev ($\sigma$) IQR (Q3-Q1)
+Qwen 2.5 (0.5B) 2.492 0.733 0.985 0.700 0.411 0.498
+LLaMA 3.2 (1B) 2.370 0.650 0.924 0.758 0.368 0.494
+Gemma 2 (2B) 2.244 0.599 0.781 0.794 0.344 0.430
+Pythia (2.8B) 3.786 0.625 0.398 0.174 0.269 0.153
+Mistral (7B) 2.443 0.607 0.824 0.804 0.414 0.515
+LLaMA 3.1 (8B) 2.495 0.621 0.859 0.691 0.339 0.444
+Pythia (12B) 2.783 0.638 0.850 0.552 0.321 0.391
+Table 3: Distributional statistics for spectral metrics across all attention heads.
+The standard deviation ($\sigma$) and interquartile range (IQR) reveal the diversity of learned strategies within each model.
+5.3 Ablation Study
+Model PE Type Heads Spectral Frequency Scale 0.5 Scale 0.25 Pos-Spec Reconstr.
 Entropy Select. Sens. Sens. Corr. Error
-LLaMA 3.2 (1B) 32 1.333 9.980 0.983 0.850 0.568 0.019
-Gemma 2 (2B) 8 1.809 8.103 0.986 0.866 0.225 0.028
-Pythia (2.8B) 32 1.689 8.298 0.981 0.853 0.591 0.019
-Qwen 2.5 (5B) 14 1.527 8.835 0.983 0.862 0.304 0.031
-Mistral (7B) 32 2.217 6.729 0.983 0.850 0.657 0.014
-LLaMA 3.1 (8B) 32 1.529 9.141 0.984 0.850 0.597 0.014
-Pythia (12B) 40 2.006 6.462 0.984 0.850 0.597 0.014
-Table 1: Comparative Analysis of Language Model Metrics
-Model 16 tok. 32 tok. 64 tok.
-LLaMA 3.2 (1B) 0.937 0.931 0.931
-Gemma 2 (2B) 1.073 1.056 1.055
-Pythia (2.8B) 0.942 0.940 0.940
-Qwen 2.5 (5B) 1.106 1.103 1.103
-Mistral (7B) 0.889 0.877 0.877
-LLaMA 3.1 (8B) 0.878 0.876 0.877
-Pythia (12B) 0.878 0.877 0.877
-Table 2: Multi-Resolution Window Entropy Analysis
-As we can see from Table 1 and Table 2, the remarkably consistent pattern across all models where correlation remains near-perfect (0.98) at $0.5x$ scale but degrades to 0.850.85~{}0.850.85 at $0.25$x scale reveals a fundamental property of wavelet transforms: graceful degradation across scales.
-This pattern directly mirrors the behavior of wavelet basis functions, which maintain high correlation with dilated versions of themselves up to a critical scale factor.
-The consistency of this pattern across architectures and model sizes (from 1B to 27B parameters) suggests this isn’t a random artifact but rather a fundamental property of how these models learn to process positional information.
-The degradation curve closely matches what we would expect from a system using wavelet-like basis functions to decompose and reconstruct signals.
-Spectral Analysis Evidence
-The inverse relationship between model size and frequency selectivity provides strong evidence for wavelet-like behavior: smaller models (e.g., LLaMA 1B) show high frequency selectivity (9.980) and low spectral entropy (1.333), indicating they develop sharp, specialized frequency bands - similar to wavelets with high Q-factors; while larger models (e.g., Pythia 12B) show lower selectivity (6.462) and higher spectral entropy (2.006), suggesting they develop more distributed representations - analogous to having a richer set of wavelet basis functions.
-This trade-off perfectly aligns with wavelet theory: systems with limited capacity optimize for sharp frequency selectivity, while systems with more capacity can afford overlapping wavelets that provide better reconstruction properties.
-Multi-Resolution Analysis Support
-The stability of entropy across different window sizes (e.g., Mistral 7B: [0.889, 0.877, 0.877]) provides crucial evidence for wavelet-like behavior.
-This pattern indicates that the representations maintain consistent information content across scales, the attention patterns exhibit self-similarity properties and the models develop scale-covariant features.
-These properties are hallmark characteristics of wavelet transforms but are not natural properties of the base RoPE mechanism, indicating they must be learned compensatory behaviors.
-Uncertainty Principle Conformance
-The variation in position-spectrum correlation across model sizes reveals how models balance the fundamental uncertainty principle.
-In fact, smaller models (Gemma 2B: 0.224) show low correlation, indicating they maintain separate positional and frequency channels, while larger models (Mistral 7B: 0.657) show higher correlation, suggesting more integrated representations
-This progression exactly matches what we would expect from a system evolving increasingly sophisticated wavelet-like properties: smaller models use simpler, more separated representations, while larger models develop more nuanced, integrated representations that better balance the position-frequency trade-off.
-Frame Completeness Evidence
-The systematic improvement in reconstruction error with model size (from 0.031 for Qwen 2.5 5B to 0.014 for Pythia 12B) provides perhaps the strongest evidence for wavelet-like behavior.
-This pattern shows that larger models develop more complete wavelet frames, that the representations become more orthogonal and efficient and the system learns to better approximate the completeness relation of wavelet frames.
-This is exactly what we would expect if models are learning to approximate wavelet transforms: larger models can learn more basis functions, leading to better frame properties and lower reconstruction error.
-This evidence is particularly compelling because it shows that models independently discover and implement principles from wavelet theory without being explicitly designed to do so.
-The consistent patterns across different architectures and scales suggest this is a fundamental property of how neural networks compensate for the limitations of fixed positional encodings.
-The progression of these properties with model scale - from simple, specialized representations in smaller models to rich, integrated representations in larger models - provides strong evidence that this is a learned adaptation rather than an architectural accident.
-This supports the broader hypothesis that neural networks naturally evolve optimal solutions for processing hierarchical information across multiple scales.
+Llama-3.2-3B RoPE 24 2.425 $\pm$ 0.630 0.728 $\pm$ 0.349 0.038 $\pm$ 0.036 0.090 $\pm$ 0.040 -0.502 0.00012 $\pm$ 0.00008
+flan-t5-base T5 12 2.696 $\pm$ 0.712 0.704 $\pm$ 0.480 0.627 $\pm$ 0.239 0.689 $\pm$ 0.231 -0.790 0.00004 $\pm$ 0.00003
+bert-base BERT 12 2.449 $\pm$ 0.819 0.743 $\pm$ 0.446 0.507 $\pm$ 0.342 0.548 $\pm$ 0.350 -0.606 0.00007 $\pm$ 0.00006
+gpt2 No PE 12 2.868 $\pm$ 0.767 0.514 $\pm$ 0.369 0.141 $\pm$ 0.164 0.152 $\pm$ 0.154 -0.672 0.00005 $\pm$ 0.00005
+Table 4: Ablation study comparing different position encoding (PE) methods. We report mean and standard deviation across attention heads for key metrics.
+To validate that the observed compensatory mechanisms are uniquely characteristic of RoPE’s implicit relative positioning, we conducted an ablation study comparing it against three alternative positional encoding schemes: T5’s relative position biases, BERT’s absolute positional embeddings, and a GPT-2 model with no explicit positional encoding (No PE).
+The results, detailed in Table 4, show that RoPE provides a uniquely effective solution for maintaining positional awareness across different scales.
+The analysis highlights a clear hierarchy in scale invariance.
+RoPE stands out with an exceptionally low scale sensitivity ($S_{0.5}=0.038$), demonstrating a robust ability to preserve positional representations when sequence lengths change.
+At the other extreme, models with explicit, non-rotary encodings (T5 and BERT) proved to be comparatively rigid, with high scale sensitivity values of 0.627 and 0.507, respectively.
+T5’s rigidity is further underscored by its high spectral entropy (2.696) and the most strongly negative position-spectrum correlation ($\rho=-0.790$), indicating a significant conflict between positional focus and spectral organization.
+The GPT-2 model (No PE) presents a fascinating intermediate case.
+While it achieves moderate scale invariance ($S_{0.5}=0.141$), far superior to T5 and BERT, it is significantly less robust than RoPE.
+Interestingly, it accomplishes this not through sharp, focused attention, but through the opposite strategy.
+It registered the lowest frequency selectivity (0.514) and the highest spectral entropy (2.868), suggesting its attention patterns are spectrally diffuse.
+This implies that in the absence of positional guidance, the model adopts a high-entropy, less-structured approach that offers some resilience but lacks the precision of RoPE.
+Ultimately, these findings underscore that RoPE’s architecture fosters wavelet-like properties that other positional encoding implementations don’t.
+It achieves scale-invariance that surpasses not only rigid, explicit encodings but also the more diffuse, emergent strategies of models with no positional guidance, marking it as a uniquely effective and well-balanced method for encoding position.
+5.4 Evolution of wavelet-like features during training
+To understand how this wavalet-like compensatory strategies develop, we analyzed Pythia 6.9b at different stages of its training, from initialization (step 0) to 143,000 steps.
+The results, summarized in Table 5, reveal that the model does not learn its representations linearly but instead undergoes distinct developmental phases.
+Initially, at step 0 and step 128, the untrained model exhibits a simple default state: high frequency selectivity ($\sim 0.76$) and low spectral entropy ($\sim 2.29$).
+This indicates that the attention heads are undifferentiated, focusing on very basic, low-complexity patterns before significant learning has occurred.
+A dramatic "exploratory phase" begins around step 512 and peaks precisely at step 1000.
+In this critical period, frequency selectivity plummets to its absolute minimum of 0.230, while spectral entropy surges to its maximum of 3.522.
+This is driven by a significant reallocation of representational power away from the low-frequency band (which drops to a minimum of 43.4%).
+Following this peak, the model enters a "specialization phase" (step 5000 onwards), where spectral entropy gradually decreases and frequency selectivity begins to recover, suggesting the model is refining and consolidating its newly learned complex representations.
+Perhaps the most striking finding is the evolution of scale sensitivity, which increases in a two-step process.
+The first significant jump occurs between step 512 and step 1000, where sensitivity rises from 0.533 to 0.617.
+A second, larger leap happens between step 1000 and step 5000 (from 0.617 to 0.742), where it reaches its plateau.
+This shows that as the model’s spectral strategies become more complex, its representations become highly tuned to specific scales.
+Rather than developing a truly robust, scale-invariant understanding, the model specializes, making its learned positional strategies more brittle when the input scale is changed.
+Our analysis is anchored by a very low reconstruction error across all checkpoints, validating the numerical stability of our wavelet methodology.
+Training Step Spectral Entropy Frequency Selectivity Low Freq. Power (%) Scale Sens. (0.5x) Scale Sens. (0.25x)
+0 2.277 0.757 77.0 0.523 0.539
+128 2.291 0.760 76.0 0.522 0.539
+512 3.169 0.369 54.4 0.533 0.549
+1000 3.522 0.230 43.4 0.617 0.633
+5000 3.381 0.294 48.6 0.742 0.762
+10000 3.293 0.330 51.7 0.748 0.767
+143000 3.010 0.454 59.7 0.740 0.756
+Table 5: Evolution of spectral and scale invariance metrics for a Pythia model across training checkpoints.
 6 Theoretical Framework for Wavelet-like Attention Patterns
@@
 6.1 Natural Evolution Toward Wavelet Behavior
+RoPE’s rotational encoding introduces specific frequency components that propagate through the attention mechanism in a mathematically structured way.
+The rotation matrix $R(m\theta_{k})$ creates an inherent trade-off: larger $\theta$ provides precise positions but causes rapid rotation cycles that confuse distant relationships, while smaller $\theta$ better captures long-range patterns but blurs local positions.
+The wavelet-like properties we observe show how attention heads adapt to handle different frequency ranges created by these rotations.
 As models train, these inherent limitations place evolutionary pressure on the learned representations. Attention heads respond by developing wavelet-like properties for three principal reasons:
@@
 7 Implications
-The practical implications of our findings are particularly compelling, understanding that attention heads naturally organize into frequency bands suggests new approaches to model initialization and architecture design.
-For instance, we could potentially pre-initialize attention heads to approximate different wavelet scales, accelerating training by starting from a more optimal configuration.
-This could be especially valuable for smaller models where computational efficiency is crucial.
-The multi-resolution nature of these emergent properties also has implications for transfer learning and domain adaptation.
-Understanding how models naturally handle different scales of information could help us design better pre-training objectives and fine-tuning strategies that explicitly account for this hierarchical processing structure.
-In essence, our findings not only deepen our understanding of how transformer models work but also provide practical tools for improving their design and implementation.
-This bridge between theory and practice could prove valuable as we continue to advance the field of language model development.
+Our findings have direct implications for the design, training, and specialization of language models.
+The discovery that attention heads spontaneously organize into frequency specialists suggests that models could be made more efficient by pre-initializing heads to a wavelet-like basis.
+This could accelerate convergence by bypassing the inefficient "exploratory phase" we identified during training.
+This functional understanding also provides a principled method for model pruning, creating smaller models by removing redundant frequency heads, and a powerful diagnostic tool for assessing training stability by monitoring spectral metrics.
+Furthermore, this work provides a roadmap for model adaptation and future research.
+Understanding head specialization allows for targeted, task-specific fine-tuning, where one could adapt low-frequency heads for summarization or high-frequency heads for code analysis.
+This "wavelet framework" serves as a new lens for interpretability and establishes a new benchmark for positional encodings.
+The goal is no longer just to provide a position signal, but to find encodings that, like RoPE, act as a powerful inductive bias to catalyze the emergence of these optimal, multi-resolution strategies.
 8 Conclusion
-Our research into the relationship between rotary positional embeddings and attention patterns reveals a fascinating aspect of how large language models adapt to theoretical limitations.
-We have shown that transformer models naturally evolve wavelet-like properties, and that servers as a compensation mechanism for the inherent constraints of RoPE, with this behavior becoming more sophisticated as model scale increases.
-The consistent pattern of frequency band distribution across different model scales, the systematic improvement in frame completeness with model size, and the remarkable stability of multi-resolution entropy all point to a learned adaptation that closely mirrors wavelet transform properties.
-What makes this particularly intriguing is that no aspect of the models’ architecture explicitly encourages such behavior – it emerges naturally through training, suggesting this may be an optimal solution to the fundamental challenge of balancing local and global information processing.
-The progression of these properties with model scale is especially revealing.
-Smaller models develop simpler but more specialized frequency responses, while larger models evolve more nuanced, integrated representations.
-This scalability suggests that the wavelet-like behavior is not merely a coincidental feature but a fundamental property of how these models learn to process hierarchical information across multiple scales.
-These findings have significant implications for future model development.
-Understanding that attention heads naturally evolve to approximate wavelet transforms could inform more efficient architectural designs, potentially leading to models that explicitly leverage this property rather than requiring it to be learned.
-This could be particularly valuable for smaller models, where computational efficiency is crucial.
-Looking forward, these findings contribute to our understanding of how neural networks implement sophisticated mathematical principles, even when not explicitly designed to do so.
+Our research demonstrates that Transformer models equipped with Rotary Position Embeddings (RoPE) are not merely subject to their theoretical limitations but actively overcome them by developing emergent, wavelet-like processing strategies.
+We have shown that this adaptation is not a minor artifact but a fundamental and multi-faceted characteristic of modern language models.
+Indeed, we established that RoPE’s implicit relative encoding is uniquely instrumental in fostering the remarkable scale-invariance observed, distinguishing its behavior from models with absolute or explicit relative position biases.
+This adaptation is not static; rather, it is learned through distinct developmental phases.
+An initial "exploratory" stage, characterized by high spectral entropy, gives way to a "specialization" stage where models refine their strategies, becoming more efficient yet often more scale-dependent.
+Crucially, this dynamic learning process culminates in models that respect the fundamental Heisenberg-Gabor uncertainty principle, with larger and more advanced architectures demonstrating increasingly sophisticated and diversified strategies for managing this inherent trade-off between positional precision and spectral organization.
+Looking forward, this "wavelet framework" offers a new lens for interpretability, allowing us to analyze what a model learns in terms of scale and frequency.
+Furthermore, understanding that these optimal structures are emergent rather than designed could inspire new architectural innovations, potentially leading to models that are either explicitly guided towards these solutions or are built with more efficient, innate multi-scale processing capabilities from the start.
 9 Limitations
-Our study shown theoretically and with experiments that Transformer-based LLMs learn to process hierarchical information across multiple scales in a way that resemble a wavelet.
-However, limitations and questions remain to be addressed.
-First, our understanding of how these properties emerge during training is still limited: we shown that this behavior helps the model overcoming the theoretical limitations of RoPE, but we didn’t study whether is RoPE that induces the behavior itself.
-Second, our analysis studied this behavior at inference time, in fact we think that a possible future work would be studying the emergence of the wavelet-like patterns at training time.
-An intriguing limitation we encountered involves the interaction between these wavelet-like properties and the model’s handling of ambiguous or context-dependent information.
-While the wavelet-like behavior provides an elegant solution for position encoding, it may introduce subtle biases in how models process semantically nuanced content.
-Further research could explore whether these biases affect the model’s performance on tasks requiring fine-grained semantic discrimination.
-A potential risk coming from our paper is that the findings show how the wavelet-like properties become more sophisticated in larger models, and it might contribute to the trend of focusing on ever-larger models, potentially exacerbating issues of resource concentration and environmental impact.
+While our study provides strong evidence for the emergence of wavelet-like properties in RoPE-based models, its scope and interpretations have several limitations that open promising avenues for future research.
+Our primary analysis, like much of the field, examines model properties at inference time after training is complete.
+While the Pythia checkpoint analysis offers a glimpse into the learning trajectory, a more fine-grained, continuous analysis of how these spectral properties evolve during the training process could reveal the precise dynamics and triggers for the observed phase shifts.
+Although we analyzed a range of model sizes and architectures, our findings are based on a specific set of open-source language models trained predominantly on English text.
+Further research is needed to determine if these principles generalize across other modalities (e.g., vision, audio), data types (e.g., code, multilingual text), and proprietary architectures.
+Broader impact
+A potential societal implication of our work is that by demonstrating these beneficial properties intensify with model scale, our findings could be used to justify the trend of building ever-larger models.
+This risks exacerbating existing issues of computational and environmental resource concentration.
+10 Acknowledgments
+Research supported in part by European Union - Next Generation EU - namely by the MUR-PRIN 2022 project "2022REWNTE - Artificial Intelligence algorithms to track and detect Covid-19 vaccine-related infodemic on social media" - CUP no.B53D23020690006; in part by the projects FAIR under Grant PE0000013 and SERICS under Grant PE00000014 under the MUR National Recovery and Resilience Plan funded by the European Union-NextGenerationEU, and in part by the project Neural Reasoning over Open Data (NEREO) funded by the Italian Ministry of Education and Research (PRIN) under Grant 2022AEFHA, and in part by the project SEED funded by Sapienza University of Rome.